OpenStudy (anonymous):

show steps to identify: (cos x + cos y)^2 + (sin x – sin y)^2 = 2 + 2cos(x + y)

OpenStudy (anonymous):

Use \[(a\pm b)^2 = a^2 \pm 2 a b + b^2 \]

OpenStudy (anonymous):

Then use \[ \cos ( x + y) =\cos (x) \cos(y) - \sin(x) \sin(y) \]

OpenStudy (anonymous):

could you please explain this a little more? I don't really understand how to use them

OpenStudy (anonymous):

Some moderators will issue warning if we try to help too much.!!!!

OpenStudy (anonymous):

\[(cos x + cos y)^2 = \cos^2(x) + 2 \cos(x) \sin(x) + \sin^2(x)= 1 + 2\cos(x) \sin(x) \]

OpenStudy (anonymous):

Do the same for the next term and expand it.

OpenStudy (callisto):

Eh... \[(cos x + cos y)^2=cos^2x + 2cosxcosy +cos^2y\]\[(sin x – sin y)^2 = sin^2x - 2sinxsiny +sin^2y\] So, \[(cos x + cos y)^2 + (sin x – sin y)^2 = (cos^2x + 2cosxcosy +cos^2y)+(sin^2x - 2sinxsiny +sin^2y)\]\[=(cos^2x +sin^2x) + (cos^2y+sin^2y) + 2(cosxcosy-sinxsiny)\]\[=1+1+ 2 cos(x+y)\]\[=can \ you \ continue?\]

OpenStudy (callisto):

For the fifth line: \[(cosx+cosy)^2 + (sinx-siny)^2\]\[=(cos^2x+2cosxcosy+cos^2y)+(sin^2x-2sinxsiny+sin^2y)\]